Math 242  University of South Carolina
D. Meade  Spring 1998

Math 242 (Section 3) – Elementary Differential Equations

Instructor Professor Doug Meade

Office Hours: TTh 2:00– 3:30 and by prior appointment
Office:  LeConte College 300E
Phone:  777-6183

 E-mail:  meade@math.sc.edu

WWW URL  http://www.math.sc.edu/~meade/math242-S98/

Meeting Times TTh 11:00am– 12:15am, LC 303a

Text \underline{Required}: Differential Equations, preliminary edition, by Blanchard, Devaney, and Hall, PWS, 1997. \underline{Recommended}: Engineer’s Toolkit: Maple V for Engineers, by Meade and Bourkoff, Addison–Wesley, 1998.

Prerequisite Completion of Math 142 with a grade of C or better.

Overview Differential equations is one of the “languages” utilized by scientists and engineers to model the physical world. In this course we will learn both how to use differential equations to describe different phenomena and how to analyze differential equations – sometimes without explicitly finding a solution.

One of the main objectives of this course is for you to understand the basic concepts of differential equations well enough to know when, how, and why to apply them in real-world situations and to be able to interpret and communicate the results.

Reaching this goal requires practice at a variety of numerical, graphical, and analytic methods. In contrast to the typical mathematics course, it is not sufficient to be proficient only with the mechanical manipulations. Instead, we will be concentrating on modeling and the understanding of differential equations. In conjuction with this approach to the material, we will learn to use state-of-the-art computer software (Maple) to perform many of the mechanical calculations.

You will also be expected to develop and practice your verbal and written communication skills. You will be required to work in groups during class. It is hoped that you will also find a group to work with outside of class.

Use of Technology The computer will be used extensively in this course. All “programming” will be done using the Maple software package. Maple provides a convenient interface to symbolic, numerical, and graphical interpretations of a variety of mathematical objects. You will be expected to become proficient in the use of Maple for the analysis of differential equations. In particular, you will need to recognize when the use of the computer is appropriate and when hand (or mental) calculations are more more efficient. Supplemental materials will be distributed at appropriate times throughout the semester.

Course Content The course will cover most of the material presented in Chapters 1–5 and 8 of the text. The text preface, and the note to the student, provide additional information about the perspective from which this course will be taught. Read these pages!

Grading Your grade in this course will be based on your performance on quizzes, two (2) midterm exams, two (2) group projects, and a final exam. The weights assigned to each of these components will be: